# Teaching Large Language Models to Maintain Contextual Faithfulness via Synthetic Tasks and Reinforcement Learning

### v1

@@ Appendix B Dual-GRPO Details
Accuracy Reward. For short-form generation, we directly assign the accuracy reward. Specifically, for the generated short-form response $y_{sf}$ based on the given context $c$ and question $q$, which is extracted from the whole generated response $y_{whole}$ via string matching, and the ground truth answer $y_{gt}$ from the synthesized training data, the accuracy reward $R_{\text{acc}}$ for the LLM $\theta$ can be calculated as: R_acc = {1 if ysf(c,q | θ) =ygt,0 otherwise.
We use the exact matching (EM) to measure accuracy, giving a score of 1 for a match and 0 for a mismatch. In this way, we can ensure that the generated short-form response correctly answers the question based on the given context, making LLMs more faithful in short-form response generation.
Proxy Reward. Instead of directly evaluating the faithfulness of the generated long-form response, we propose a proxy reward to evaluate it implicitly. Specifically, for each generated long-form answer $y_{lf}$, we replace the given context $c$ with it as new input and infer the LLM $\theta$ to determine whether the LLM can produce the correct short-form answer $y_{sf}$ based on $y_{lf}$ for the question $q$. Thus, the proxy reward $R_{\text{proxy}}$ can be calculated as: R_proxy = {1 if ysf(ylf,q| θ) =ygt,0 otherwise.
If the generated long-form response can help LLMs generate the correct answer, it indicates that the long-form response is faithful to the context, contains syntactically and semantically complete sentences, and correctly addresses the question. Thus, we assign a reward score of 1 for the positive long-form response that helps the LLM to produce the correct answer, and a reward score of 0 for those that lead to incorrect answers.
Format Reward. To enforce the desired output format, we assign a reward on the whole generated response $y_{whole}$ to evaluate whether it contains the proper XML tags. We use three types of tags as shown in our system prompt, as shown in Figure 11, including <think>, <long_answer>, and <short_answer> tags. Formally, R_format = {1 if correct formatting is present,0 if incorrect formatting.
@@
Final Reward. Finally, we use the sum of these three rewards as the final composite reward $R_{\text{final}}$. This well-designed reward $R_{\text{final}}$ of Dual-GRPO enhances the efficacy of the rule-based RL training framework to guide the model toward generating more faithful responses in both short-form and long-form tasks. Formally, R_final = R_acc + R_proxy + R_format.
Finally, we use this reward $R_{\text{final}}$ to compute an advantage $A_{i}$ for each output, guiding policy updates according to the GRPO objective.
@@ C.1 Datasets
CNQ (Counterfactual QA). CNQ is constructed based on Natural Questions (Kwiatkowski et al., 2019a). In CNQ, the context is modified to support counterfactual answers following (Longpre et al., 2021). It contains 2,773 samples that incorporate counterfactual passages to evaluate the faithfulness of LLMs on short-form generation.
FaithEval (Counterfactual Multiple-choice QA). FaithEval is a novel and comprehensive benchmark tailored to evaluate the faithfulness of LLMs in contextual scenarios across three diverse tasks: unanswerable, inconsistent, and counterfactual contexts. We select the counterfactual task to evaluate the faithfulness of LLMs, which contains 1,000 multiple-choice QA samples curated based on ARC-Challenge (Clark et al., 2018).
FiQA (Factual QA). FiQA is a factual version of ConFiQA, which shares the same questions as ConFiQA but contains the factual contexts and answers. The contexts and answers are provided by Bi et al. (2024), thus we can evaluate the faithfulness of LLMs in factual short-form response generation. It contains 1,500 samples for evaluation.
FollowRAG (RAG Scenarios for short-form QA). FollowRAG aims to assess the model’s ability to follow user instructions in complex multi-document contexts. It consists of four well-known open-domain QA datasets for RAG scenarios, including NaturalQA, TriviaQA, HotpotQA, and WebQSP. We utilize the provided passages in FollowRAG as context and original query (instead of the version with added instruction constraints proposed by Dong et al. (2024)) as questions. We also use the original answers to report the results. FollowRAG contains 2,800 samples used for testing (700/700/700/700 from NaturalQA/TriviaQA/HotpotQA/WebQSP). Different from short-form generation tasks that the contexts always contain answers, in real-world RAG scenarios, the answer may not appear in the retrieved passages, and these passages tend to be noisy.
XSum (Summarization). Summarization is a content-grounded task where a model is provided a piece of text and tasked with synthesizing the most salient information within that text. XSum is a widely used dataset for text summarization, which consists of about 220,000 BBC articles as input documents. To facilitate our evaluation, we use the first 1,000 data points from the test set to evaluate our method.
WikiLarge (Simplification). Text simplification is a content-grounded task where a model is provided a piece of text and is tasked with paraphrasing it to make the text easier to read and understand. We use 1k instances sampled from the WikiLarge dataset as a test set, following Ravichander et al. (2025).
@@ C.2 Metrics and LLM-as-a-Judge
Metrics for Short-form Generation Tasks. We evaluate performance based on whether gold answers are included in the generated responses (i.e., Acc) following Asai et al. (2024) and exact matching (EM) for QA tasks. For multiple-choice questions in FaithEval, we use keyword matching to verify the accuracy, i.e., Acc.
Metrics for Long-form Generation Tasks. To evaluate the faithfulness of generated long-form answers, we use MiniCheck to check whether the model response is grounded in the provided context. MiniCheck is a state-of-the-art method to recognize if LLM output can be grounded in given contexts. We select the MiniCheck-FT5222https://huggingface.co/lytang/MiniCheck-Flan-T5-Large because it is the best fact-checking model, outperforming GPT-4o in evaluating the faithfulness. If the model response contains at least one statement that cannot be inferred from the context, we consider it as a negative response; otherwise, it is a positive response. To evaluate the quality of the generated long-form responses for three different tasks (QualityScore), including summarization, simplification, and long-form QA, we design different prompts to query GPT-4o-2024-11-20 as a judge to get the quality scores. We report the average results of the quality score results by querying GPT-4o twice. The prompts for three tasks can be found in Figure 12, Figure 13, and Figure 14.
@@ C.3 Baselines
C.4 Test-time Prompts
For baselines, the prompts for different tasks can be found in Figure 15, Figure 16, Figure 17, Figure 18, and Figure 19. To evaluate the factuality of LLMs, we modify the original FaithEval and make it a closed-book QA setting, and use the prompts shown in Figure 20. During the evaluation for Canoe, we apply the same system prompt during the Dual-GRPO training, and extract the content between <short_answer> and </short_answer> tags as the final answers for short-form generation tasks. Also, for long-form generation tasks, we extract the content between <long_answer> and </long_answer> tags as the final answers. We also find that the long-form responses generated by Canoe can provide correct answers in short-form generation tasks in the Appendix F.1. Thus, for real-world applications, we recommend using the generated long-form responses as the system responses for the user’s instructions, because these long-form responses can not only faithfully complete long-form generation tasks, but also provide correct answers in short-form generation tasks.
@@ C.4 Test-time Prompts
[…]
Table 7: Experimental results (%) on FollowRAG. Bold numbers indicate the best performance of models with the same model size.
C.5 More Detailed Experimental Results
@@
Appendix D Implementations Details
We implement our method based on the RL framework open-r1 (Face, 2025). We use AdamW optimizer (Loshchilov and Hutter, 2019) to train our model, with a $1\times 10^{-6}$ learning rate, a batch size of 14 for 7B/8B models, and a batch size of 7 for the 14B model, steering the training across two epochs. We set the maximum input length for the models to 1,024 and the maximum generation length to 1,024. The number of generations $G$ during the RL training is set to 7, which is used in Eq. (1). We set 0.04 for $\beta$ used in Eq. (1). We set 0.2 for $\epsilon$ used for the clip shown in Eq. (2). We set 0.9 for temperature in RL training to generate responses. We conduct our experiments on NVIDIA A800-80G GPUs with DeepSpeed+ZeRO2 for 7B/8B models, DeepSpeed+ZeRO2+Offloading for the 14B model, and BF16. During the inference, we set 0.7 for temperature for the evaluation of our models and baselines. For each task, we infer the model twice and report the average scores as final results.
Appendix E Human Evaluation
We conduct a human evaluation on the 90 samples from long-form generation tasks, including 30/30/30 for summarization/simplification/long-form QA. We evaluate these samples across four key dimensions: readability, faithfulness, helpfulness, and naturalness. For each comparison, three options are given (Ours Wins, Tie, and Initial Model Wins), and the majority voting determines the final result. The participants follow the principles in Figure 21 to make the decision. We invite three Ph.D. students to compare the responses generated by the models. Before participants begin to make judgments, we describe the principles of our design in detail and ensure that each participant correctly understands the principles. If the final result can not be determined by majority voting, we will hold a discussion among the participants and vote on the result again. We compare two models, including Canoe-LLaMA-8B as our method and LLaMA-3-8B as the initial model. Shown in Figure 5, we can find that our method reduces faithfulness hallucinations and also ensures the response quality for three long-form generation tasks.
@@ Appendix F Discussion
F.1 Can Long-form Responses Generated by Canoe Provide Correct Answers in Short-form Generation Tasks?
This exploration is important because, in real-world applications, it is difficult to pre-determine whether to use generated short-form responses (i.e., the context between <short_answer> and </short_answer> tags) or long-form responses (i.e., the context between <long_answer> and </long_answer> tags) as answers to respond to user instructions. This contrasts with the evaluation of LLMs on different datasets, as described in the test-time strategies outlined in C.4. Therefore, we first explore whether the long-form responses generated by Canoe (i.e., the context between <long_answer> and </long_answer> tags) can provide correct answers in short-form generation tasks. As shown in Table 8, when evaluating the generated long-form responses that contain the free-form answers, the accuracy scores consistently increase in all the short-form generation tasks compared to using the generated short-form responses. It also indicates that the generated short-form responses maintain conciseness, which is important for measuring the EM score, but can slightly reduce the accuracy score. Therefore, in real-world applications, we can directly use the generated long-form responses as the system responses for the user’s instructions, because these long-form responses can not only efficiently and faithfully complete long-form generation tasks, but also provide correct answers in short-form generation tasks.
Model Accuracy Proxy Format
Canoe-LLaMA-8B 70.3 66.1 99.4
Canoe-Qwen-7B 64.1 63.4 99.9
Canoe-Qwen-14B 83.5 76.5 100.0
@@ F.1 Can Long-form Responses Generated by Canoe Provide Correct Answers in Short-form Generation Tasks?
F.2 Final Rewards in the RL Training Stage
We show the final rewards in Table 9. We can find that models can easily learn the designed format, while accuracy and proxy rewards still remain challenging. Meanwhile, in the early stages of RL training, the format reward increases quickly and converges rapidly, and as training proceeds, the accuracy reward and the proxy reward gradually increase and eventually converge. This indicates that our well-designed training data construction strategy is effective and ensures the complexity and diversity, avoiding overfitting and reward hacking.
@@ F.2 Final Rewards in the RL Training Stage
F.3 Multilingual Transfer Ability and Context Length Generalization of Canoe
To further explore the multilingual transfer ability of Canoe, we further evaluate our model on the Chinese dataset. Specifically, we use the single-document QA dataset MultiFieldQA-zh (Bai et al., 2023), the multi-document QA dataset DuReader (He et al., 2018), and the summarization dataset VCSUM (Wu et al., 2023) within LongBench (Bai et al., 2023). Following Si et al. (2024) that utilizes the GPT-4 to evaluate the correctness of QA tasks and the faithfulness of the summarization task, we use the advanced LLM to evaluate these datasets. We use the same prompts for three tasks as Si et al. (2024) to query GPT-4o. The test-time prompts for these tasks can be found in Figure 23, Figure 24, and Figure 25. As shown in Table 10, we can find that our proposed Canoe also improves the faithfulness in Chinese datasets, indicating that our proposed method has a strong multilingual transfer ability. Meanwhile, these results also indicate that our method achieves better faithfulness even when our training data is short, i.e., Canoe guarantees a consistently strong performance of the LLM in long-context scenarios.
Figure 6: The Avg EM results (%) on 11 datasets with different numbers of synthesized short-form training data. We conduct the experiments based on LLaMA-3-Instruct-8B models.
F.4 Effect of the Amount of the Synthesized Short-form Data
To further explore the effect of the amount of the synthesized short-form data, we conduct the corresponding experiments as shown in Figure 6. We can find that increasing the number of training data improves the performance of our method, but when the number of data is greater than 10,000, the performance will be stable. For the models with different sizes, the optimal amount of training data may be different, e.g., the larger models may need more training data to achieve optimal performance. Fortunately, due to our training data construction strategy, we can simply scale and synthesize training data without human annotation.
@@ F.4 Effect of the Amount of the Synthesized Short-form Data
[…]
@@
[…]
@@
[…]
@@
F.5 Case Study between GRPO and the proposed Dual-GRPO
We find that directly applying GRPO instead of our proposed Dual-GRPO to synthesized short-form data leads to over-optimizing short-form generation and a false response generation pattern. The used system prompt for applying GRPO can be found in Figure 22. Shown in Table F.4, we can find that the tuned model GRPO-LLaMA-8B tends to directly copy text spans from the given context as the final answer instead of following instructions in long-form generation tasks. However, when we apply Dual-GPRO to our synthesized data, we find that trained models can generate fluent and complete sentences. Thus, Dual-GRPO not only improves the faithfulness of LLMs in two types of response generation but also ensures the utility of models.
@@ F.5 Case Study between GRPO and the proposed Dual-GRPO
Table 12: Manually selected relations that are used to construct training data. We utilize the same manually selected relations as Bi et al. (2024).
Prompt for question generation for the samples with straightforward context. [Instructions] You are a sophisticated question generator. Given a triple {$(h,r,t)$} collected from Wikidata, generate a question that asks about the final tail entity {$t$} using the head entity {$h$} and the relation {$r$}. Directly give me the generated question:
@@
Prompt for context generation for the samples with straightforward context. [Instructions] You are a sophisticated context generator. Given a triple {$(h,r,t)$} collected from Wikidata, generate a brief description of the head entity {$h$}, approximately 150 words long. Ensure the tail entity {$t$} and relation {$r$} are accurately mentioned in the generated description. Directly give me the generated context:
@@
Prompt for question generation for the samples with reasoning-required context. [Instructions] You are a sophisticated question generator. Given a chain of triples {[…]} collected from Wikidata, generate a question that asks about the final tail entity {$t$} using the head entity {$h$} and the relation {$r$}. Do not include any bridge entities in the question; instead, phrase the question as if directly asking about the relationship from the head entity to the tail entity Directly give me the generated question:
@@
Prompt for context generation for the samples with reasoning-required context. [Instructions] You are a sophisticated context generator. Given a chain of triples {[…]} collected from Wikidata, generate a brief description of the head entity {$h$}, approximately {150*n} words long. Ensure the tail entity {$t$} and relation {$r$} are accurately mentioned in the generated description. Directly give me the generated context:
@@
Test-time prompt used for MultiField-zh. {CJK} UTF8gbsn 阅读以下文字并用中文简短回答:{} 现在请基于上面的文章回答下面的问题,只告诉我答案,不要输出任何其他字词。 问题:{} 回答:
@@
Test-time prompt used for DuReader. {CJK} UTF8gbsn 请基于给定的文章回答下述问题。 文章:{} 问题:{} 回答:
@@
Test-time prompt used for VCSUM. {CJK} UTF8gbsn 下面有一段会议记录,请你阅读后,写一段总结,总结会议的内容。 会议记录:{} 会议总结:

### v3

@@ Appendix B Dual-GRPO Details
Accuracy Reward. For short-form generation, we directly assign the accuracy reward. Specifically, for the generated short-form response $y_{sf}$ based on the given context $c$ and question $q$, which is extracted from the whole generated response $y_{whole}$ via string matching, and the ground truth answer $y_{gt}$ from the synthesized training data, the accuracy reward $R_{\text{acc}}$ for the LLM $\theta$ can be calculated as:
$R_{\text{acc}}=\begin{cases}1&\text{if }y_{sf}(c,q|\theta)=y_{gt},\\ 0&\text{otherwise}.\end{cases}$
We use the exact matching (EM) to measure accuracy, giving a score of 1 for a match and 0 for a mismatch. Thus, we can ensure that the generated short-form response correctly answers the question based on the given context, making LLMs more faithful in short-form response generation.
Proxy Reward. Instead of directly evaluating the faithfulness of the generated long-form response, we propose a proxy reward to evaluate it implicitly. For each generated long-form answer $y_{lf}$, we replace the given context $c$ with it as new input and infer the LLM $\theta$ to determine whether the LLM can produce the correct short-form answer $y_{sf}$ based on $y_{lf}$ for the question $q$. Proxy reward $R_{\text{proxy}}$ can be calculated as:
$R_{\text{proxy}}=\begin{cases}1&\text{if }y_{sf}(y_{lf},q|\theta)=y_{gt},\\ 0&\text{otherwise}.\end{cases}$
If the generated long-form response can help LLMs generate the correct answer, it indicates that the long-form response is faithful to the context, contains meaningful sentences, and correctly addresses the question. Thus, we assign a reward score of 1 for the positive long-form response that helps the LLM to produce the correct answer, and a reward score of 0 for those that lead to incorrect answers.
Format Reward. To enforce the desired output format, we assign a reward on the whole generated response $y_{whole}$ to evaluate whether it contains the proper XML tags. We use three types of tags as shown in our system prompt, as shown in Figure 11, including <think>, <long_answer>, and <short_answer>tags. Formally,
$R_{\text{format}}=\begin{cases}1&\text{if correct formatting is present,}\\ 0&\text{if incorrect formatting}.\end{cases}$
@@
Final Reward. Finally, we use the sum of these three rewards as the final composite reward $R_{\text{final}}$. This well-designed reward $R_{\text{final}}$ of Dual-GRPO enhances the efficacy of the rule-based RL training framework to guide the model toward generating more faithful responses in both short-form and long-form tasks. Formally,
$R_{\text{final}}=R_{\text{acc}}+R_{\text{proxy}}+R_{\text{format}}.$
Finally, we use this reward $R_{\text{final}}$ to compute an advantage $A_{i}$ for each output, guiding policy updates according to the rule-based GRPO objective.
@@ C.1 Datasets
CNQ (Counterfactual QA). CNQ is constructed based on Natural Questions (kwiatkowski-etal-2019-natural). In CNQ, the context is modified to support counterfactual answers following (longpre-etal-2021-entity). It contains 2,773 samples that incorporate counterfactual passages to evaluate the faithfulness of LLMs on short-form generation.
FaithEval (Counterfactual Multiple-choice QA). FaithEval is a novel and comprehensive benchmark tailored to evaluate the faithfulness of LLMs in contextual scenarios across three diverse tasks: unanswerable, inconsistent, and counterfactual contexts. We select the counterfactual task to evaluate the faithfulness, which contains 1,000 multiple-choice QA samples curated based on ARC-Challenge (clark2018thinksolvedquestionanswering).
FiQA (Factual QA). FiQA is a factual version of ConFiQA, which shares the same questions as ConFiQA but contains the factual contexts and answers. The contexts and answers are provided by bi2024contextdpoaligninglanguagemodels, thus we can evaluate the faithfulness of LLMs in factual short-form response generation. It contains 1,500 samples for evaluation.
FollowRAG (RAG Scenarios for short-form QA). FollowRAG aims to assess the model’s ability to follow user instructions in complex multi-document contexts. It consists of four well-known open-domain QA datasets for RAG scenarios, including NaturalQA, TriviaQA, HotpotQA, and WebQSP. We utilize the provided passages in FollowRAG as context and original query (instead of the version with added instruction constraints proposed by dong2024generalinstructionfollowingalignmentretrievalaugmented) as questions. We also use the original answers to report the results. FollowRAG contains 2,800 samples used for testing (700/700/700/700 from NaturalQA/TriviaQA/HotpotQA/WebQSP). Different from short-form generation tasks that the contexts always contain answers, in real-world RAG scenarios, the answer may not appear in the retrieved passages, and these passages tend to be noisy.
XSum (Summarization). Text summarization is a content-grounded task where a model is provided a piece of text and tasked with synthesizing the most salient information within that text. XSum is a widely used dataset for text summarization, which consists of about 220,000 BBC articles as input documents. To facilitate our evaluation, we use the first 1,000 data points from the test set to evaluate our method.
WikiLarge (Simplification). Simplification is a task where a model is provided a piece of text and is tasked with paraphrasing it to make the text easier to read and understand. We use 1k instances sampled from the WikiLarge dataset as a test set, following ravichander2025halogenfantasticllmhallucinations.
@@ C.2 Metrics and LLM-as-a-Judge
Metrics for Short-form Generation Tasks. We evaluate performance based on whether gold answers are included in the generated responses (i.e., Acc) following asai2024selfrag and exact matching (EM) for QA tasks. For multiple-choice questions in FaithEval, we use keyword matching to verify the accuracy, i.e., Acc.
Metrics for Long-form Generation Tasks. To evaluate the faithfulness in long-form generation tasks, we use MiniCheck to check whether the model response is grounded in the provided context. MiniCheck is the SOTA method to recognize if LLM output can be grounded in given contexts. We select the MiniCheck-FT5222https://huggingface.co/lytang/MiniCheck-Flan-T5-Large because it is the best fact-checking model, outperforming GPT-4o in evaluating the faithfulness. If the model response contains at least one statement that cannot be inferred from the context, we consider it as a negative response; otherwise, it is a positive response. To evaluate the quality of the generated long-form responses for three different tasks (QualityScore), including summarization, simplification, and long-form QA, we design different prompts to query GPT-4o-2024-11-20 as a judge to get the quality scores. We report the average results of the quality score results by querying GPT-4o twice. The prompts for three tasks can be found in Figure 12, Figure 13, and Figure 14.
@@ C.3 Baselines
[2pt/3pt] Canoe-Qwen-14B 19.9 25.7 41.9 51.6 63.3 71.7 59.4 69.3
Table 8: Experimental results (%) on FollowRAG. Bold numbers indicate the best performance of models with the same size.
C.4 More Detailed Experimental Results
@@
C.5 Test-time Prompts
For baselines, the prompts for different tasks can be found in Figure 15, Figure 16, Figure 17, Figure 18, and Figure 19. It is worth noting that some test-time prompt templates differ from those in the original paper, which may lead to some variation in the reported results. For example, in the original paper of Context-DPO (bi2024contextdpoaligninglanguagemodels), the authors used a very simple prompt template, i.e., {Passages} Q: {} A:, to report performance results of popular LLMs on the ConFiQA dataset. Such a simple prompt template will lead to a very low EM score when the grounded-truth answer is just an entity, which may lead to unfair comparison and useless evaluation. Therefore, we redesign the corresponding prompt templates by adding length constraints as shown in Figure 15 to ensure a fair comparison. Meanwhile, since the authors of FaithEval (ming2025faitheval) do not provide official evaluation scripts and a specific task prompt for FaithEval-Counterfactual, we write our evaluation scripts and use the prompt shown in Figure 16 in our experiments. We find that the state-of-the-art LLMs, e.g., GPT-4o, show stable performance consistent with the original papers. At the same time, we observe that slightly smaller models, such as LLaMA-3-Instruct-8B, show decreased performance on our designed prompt templates. To evaluate the factuality of LLMs, we modify the original FaithEval and make it a closed-book QA setting, and use the prompts shown in Figure 20.
During the evaluation for Canoe, we apply the same system prompt during the Dual-GRPO training, and extract the content between <short_answer>and </short_answer>tags as the final answers for short-form generation tasks. Also, for long-form generation tasks, we extract the content between <long_answer>and </long_answer>tags as the final answers. We also find that the long-form responses generated by Canoe can provide correct answers in short-form generation tasks in the Appendix F. Thus, for real-world applications, we recommend using the generated long-form responses as the system responses for the user’s instructions, because these long-form responses can not only faithfully complete long-form generation tasks, but also provide correct answers in short-form generation tasks.
@@ C.5 Test-time Prompts
Appendix D Implementation Details
We implement our method based on the RL framework open-r1 (openr1). We use AdamW optimizer (loshchilov2019decoupledweightdecayregularization), with a $1\times 10^{-6}$ learning rate, a batch size of 14 for 7B/8B models, and a batch size of 7 for the 14B model, steering the training across two epochs. We set the maximum input length for the models to 1,024 and the maximum generation length to 1,024. The number of generations $G$ during the RL training is set to 7, which is used in Eq. (1). We set 0.04 for $\beta$ used in Eq. (1). We set 0.2 for $\epsilon$ used for the clip shown in Eq. (2). We set 0.9 for temperature in RL training to generate responses. We conduct our experiments on NVIDIA A800-80G GPUs with DeepSpeed+ZeRO2 for 7B/8B models, DeepSpeed+ZeRO2+Offloading for the 14B model, and BF16. During the inference, we set 0.7 for temperature for the evaluation of our models and baselines. For each task, we infer twice and report the average scores as final results.
@@ Appendix D Implementation Details
Appendix E Human Evaluation
Evaluating long-form generation tasks remains challenging. Thus, we conduct a human evaluation on the 90 samples from long-form generation tasks, including 30/30/30 for summarization/simplification/long-form QA tasks. We evaluate these samples across four key dimensions: readability, faithfulness, helpfulness, and naturalness. For each comparison, three options are given (Ours Wins, Tie, and Initial Model Wins), and the majority voting determines the final result. The participants follow the principles in Figure 21 to make the decision. We invite three Ph.D. students to compare the responses generated by the models. Before participants begin to make judgments, we describe the principles of our design in detail and ensure that each participant correctly understands the principles. If the final result can not be determined by majority voting, we will hold a discussion among the participants and vote on the result again. We compare two models, including Canoe-LLaMA-8B as our method and LLaMA-3-Instruct-8B as the initial model. As shown in Figure 5, we can find that Canoe reduces faithfulness hallucinations and also ensures the response quality for three long-form generation tasks.
@@ Appendix F Discussion
Can Long-form Responses Generated by Canoe Provide Correct Answers in Short-form Generation Tasks? This exploration is important because, in real-world applications, it is difficult to pre-determine whether to use generated short-form responses (i.e., the context between <short_answer>and </short_answer>tags) or long-form responses (i.e., the context between <long_answer>and </long_answer>tags) as answers to respond to user instructions. This contrasts with the evaluation of LLMs on different datasets, as described in the test-time strategies outlined in C.5. Thus, we first explore whether the long-form responses generated by Canoe (i.e., the context between <long_answer>and </long_answer>tags) can provide correct answers in short-form generation tasks. As shown in Table 9, when evaluating the generated long-form responses that contain the free-form answers, the accuracy scores consistently increase in all the short-form generation tasks compared to using the generated short-form responses. It indicates that the generated short-form responses maintain conciseness, which is important for measuring the EM score, but can slightly reduce the accuracy score. Therefore, in real-world applications, we can directly use the generated long-form responses as the system responses for the user’s instructions, because these long-form responses can not only efficiently and faithfully complete long-form generation tasks, but also provide correct answers in short-form generation tasks.
black Model Accuracy Proxy Format Canoe-LLaMA-8B 70.3 66.1 99.4 Canoe-Qwen-7B 64.1 63.4 99.9 Canoe-Qwen-14B 83.5 76.5 100.0 black
@@
Final Rewards in the RL Training Stage. We show the final rewards in Table 10. We can find that models can easily learn the designed format, while accuracy and proxy rewards still remain challenging. Meanwhile, in the early stages of RL training, the format reward increases quickly and converges rapidly, and as training proceeds, the accuracy reward and the proxy reward gradually increase and eventually converge. This indicates that our well-designed training data construction strategy is effective and ensures the complexity and diversity, avoiding overfitting and reward hacking.
@@
Figure 6: The Avg EM results (%) on 11 datasets with different numbers of synthesized training data. We conduct the experiments based on LLaMA-3-Instruct-8B models.
Multilingual Transfer Ability and Context Length Generalization of Canoe. To further explore the multilingual transfer ability of Canoe, we further evaluate our model on the Chinese dataset. Specifically, we use the single-document QA dataset MultiFieldQA-zh (bai2023longbench), the multi-document QA dataset DuReader (he-etal-2018-dureader), and the summarization dataset VCSUM (wu-etal-2023-vcsum) within LongBench (bai2023longbench). Following si-etal-2025-gateau that utilizes the GPT-4 to evaluate the correctness of QA tasks and the faithfulness of the summarization task, we use the advanced LLM to evaluate these datasets. We use the same prompts for three tasks as si-etal-2025-gateau to query GPT-4o. The test-time prompts for these tasks can be found in Figure 23, Figure 24, and Figure 25. As shown in Table 11, we can find that our proposed Canoe also improves the faithfulness in Chinese datasets, indicating that our proposed method has a strong multilingual transfer ability. Meanwhile, these results also indicate that our method achieves better faithfulness even when our training data is short, i.e., Canoe guarantees a consistently strong performance in long-context scenarios.
Effect of the Amount of the Synthesized Short-form Data. To further explore the effect of the amount of the synthesized short-form data, we conduct the corresponding experiments as shown in Figure 6. We can find that increasing the number of training data improves the performance of our method, but when the number of data is greater than 10,000, the performance will be stable. For the models with different sizes, the optimal amount of training data may be different, e.g., the larger models may need more training data to achieve optimal performance. Fortunately, due to our training data construction strategy, we can simply scale and synthesize training data without human annotation.
@@
[0.8pt/2pt] GRPO-LLaMA-8B: Ballet. ()
[0.8pt/2pt] Canoe-LLaMA-8B: Based on the passage, ballet is a type of dance that developed from lavish court spectacles in Italy and France. It is a combination of music, drama, poetry, song, costumes, and dance, and involves a combination of different elements, including music, costumes, and performance. The passage also mentions that members of the court nobility took part as performers, and that professional dancers began to take the place of court amateurs. This suggests that ballet is a type of dance that is characterized by its combination of different art forms and its performance by trained dancers. ()
@@
[0.8pt/2pt] GRPO-LLaMA-8B: Pocket Monsters Red and Green. ()
[0.8pt/2pt] Canoe-LLaMA-8B: The Pokémon series started with the release of Pokémon Red and Green in Japan, which were later enhanced and reprogrammed for international release. ()
@@
[0.8pt/2pt] GRPO-LLaMA-8B: Comedy clubs. ()
[0.8pt/2pt] Canoe-LLaMA-8B: Kareem Badr and his friends were able to turn around the struggling comedy club, the Hideout, by applying their skills in problem-solving and efficiency. They broke down the business into small chunks and figured out how to make each aspect better and more efficient. This approach helped them to improve the club’s performance and eventually expand it. The passage also mentions that the growing popularity of stand-up comedy is fueled in part by younger audiences, who view humorous TV hosts as their primary source for news. Many comedy clubs now require talent, and there is a career and a living to be made by good comedians who are not superstars. The clubs also make money from selling drinks and dinners, as well as from cover charges and talent fees. In addition, digital media has given live comedy a boost, with one terrific joke being able to get a million hits online. ()
@@
Case Study between GRPO and the proposed Dual-GRPO. We find that directly applying GRPO instead of our proposed Dual-GRPO to synthesized short-form data leads to over-optimizing short-form generation and a false response generation pattern. The used system prompt for applying GRPO can be found in Figure 22. Shown in Table F, we can find that the tuned model GRPO-LLaMA-8B tends to directly copy text spans from the given context as the final answer instead of following instructions in long-form generation tasks. However, when we apply Dual-GPRO to our synthesized data, we find that trained models can generate fluent and complete sentences. Thus, Dual-GRPO not only improves the faithfulness of LLMs in two types of response generation but also ensures the utility of models.
@@
Table 13: Manually selected relations that are used to construct training data. We utilize the same relations as bi2024contextdpoaligninglanguagemodels.
Prompt for question generation for the samples with straightforward context. [Instructions] You are a sophisticated question generator. Given a triple {(h,r,t)(h,r,t)} collected from Wikidata, generate a question that asks about the final tail entity {tt} using the head entity {hh} and the relation {rr}. Directly give me the generated question:
@@
Prompt for context generation for the samples with straightforward context. [Instructions] You are a sophisticated context generator. Given a triple {(h,r,t)(h,r,t)} collected from Wikidata, generate a brief description of the head entity {hh}, approximately 150 words long. Ensure the tail entity {tt} and relation {rr} are accurately mentioned in the generated description. Directly give me the generated context:
@@
Prompt for question generation for the samples with reasoning-required context. [Instructions] You are a sophisticated question generator. Given a chain of triples {[…]} collected from Wikidata, generate a question that asks about the final tail entity {tt} using the head entity {hh} and the relation {rr}. Do not include any bridge entities in the question; instead, phrase the question as if directly asking about the relationship from the head entity to the tail entity Directly give me the generated question:
@@
Prompt for context generation for the samples with reasoning-required context. [Instructions] You are a sophisticated context generator. Given a chain of triples {[…]} collected from Wikidata, generate a brief description of the head entity {hh}, approximately {150*n} words long. Ensure the tail entity {tt} and relation {rr} are accurately mentioned in the generated description. Directly give me the generated context:
@@
Test-time prompt used for MultiField-zh. 阅读以下文字并用中文简短回答:{} 现在请基于上面的文章回答下面的问题,只告诉我答案,不要输出任何其他字词。 问题:{} 回答:
@@
Test-time prompt used for DuReader. 请基于给定的文章回答下述问题。 文章:{} 问题:{} 回答:
@@
Test-time prompt used for VCSUM. 下面有一段会议记录,请你阅读后,写一段总结,总结会议的内容。 会议记录:{} 会议总结: